AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Jan 19 11:17:50 2024
Author
*
:
Subject
*
:
12.20 CARME IE514 4.1e-11mbar, ambient temperature 16.7 deg C FEE64 cabling complete WR source - VETAR2 borrowed from Nik Kurz and opwerating standalone ( no BuTIS reference ) All MACBs mode 0x3 15.20 Power ON FEE64s #1 - #15 aida02 did not boot - /etc/dhcp/dhcpd.conf typo - fixed aida04 did not boot - AIDA FEE64 PSU output #4 fuses open circuit - fuses replaced from AIDA FEE64 PSU spare channel ( aida16 not connected yet ) - fixed aida07 did not boot - power OK - FEE64 comfort LEDs OK - network switch link LED OK - network config? aida08 boots - power OK - comfort LEDs OK - network switch link LED OK - requests RFS / but does not request RFS MIDAS - can ping aida08 OK, cannot telnet aida08 'connection refused' aida09 did not boot - FEE64 comfort LEDS OFF - swap FEE64 #9 & #10 power cables - aida09 boots OK, aida10 does not boot Restore #9 & #10 power cables - swap AIDA FEE64 outputs #1 & #2 ( aida09 & aida10 ) - aida09 boots OK Conclude AIDA FEE64 PSU ( no label ) output #1 is faulty. Replace AIDA FEE64 PSU ( label ISOL3 ). aida09 & aida10 boot OK aida15/aida16 MAC addresses were swapped in https://elog.ph.ed.ac.uk/CARME/467 - fixed aida01-aida06, aida09-aida15 boot OK - aida07 does not boot, aida08 does not complete boot 16.25 Configure MIDAS DEE64 servers aida01 - aida06, aida10 - aida15 FEE64 temperatures - attachment 1 aida06 looks low ( c. 25 deg C cf. c. 40 deg C ) System wide checks - attachments 2-8 aida02 fails WR decoder WR timestamps ( attachments 7 & 8 ) look odd/low - values always c. 0x2??? 17.30 CARME IE514 OFF WR NIM emulator SMA cables to root MACB ( rear panel ) disconnected at front panel Power cycle FEE64s System wide checks & WR timestamp control - attachments 9-14 no change 17.45 Detector bias -20, -100 & 150V - attachments 15-17 CAEN N1419ET non-floating outputs ( to be checked ), HV cable SHV - 2x Lemo 00.250, LK1 not fitted 18.40 Detector bias -150V attachment 18 To Do Check ASIC settings Configure HV Check LKs Ground cabling
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
Screenshot_from_2024-01-19_16-30-27.png
Original size: 1479x1014
Attachment 2:
Screenshot_from_2024-01-19_16-23-24.png
Original size: 1479x1014
Attachment 3:
Screenshot_from_2024-01-19_16-23-33.png
Original size: 1479x1014
Attachment 4:
Screenshot_from_2024-01-19_16-23-51.png
Original size: 1479x1014
Attachment 5:
Screenshot_from_2024-01-19_16-24-02.png
Original size: 1479x1014
Attachment 6:
Screenshot_from_2024-01-19_16-24-07.png
Original size: 1479x1014
Attachment 7:
Screenshot_from_2024-01-19_16-24-44.png
Original size: 1479x1014
Attachment 8:
Screenshot_from_2024-01-19_16-31-32.png
Original size: 1479x1014
Attachment 9:
Screenshot_from_2024-01-19_17-48-30.png
Original size: 1479x1014
Attachment 10:
Screenshot_from_2024-01-19_17-48-38.png
Original size: 1479x1014
Attachment 11:
Screenshot_from_2024-01-19_17-48-46.png
Original size: 1479x1014
Attachment 12:
Screenshot_from_2024-01-19_17-48-51.png
Original size: 1479x1014
Attachment 13:
Screenshot_from_2024-01-19_17-49-11.png
Original size: 1479x1014
Attachment 14:
Screenshot_from_2024-01-19_17-49-23.png
Original size: 1479x1014
Attachment 15:
Screenshot_from_2024-01-19_17-43-55.png
Original size: 734x499
Attachment 16:
Screenshot_from_2024-01-19_17-54-33.png
Original size: 734x499
Attachment 17:
Screenshot_from_2024-01-19_17-56-01.png
Original size: 734x499
Attachment 18:
Screenshot_from_2024-01-19_18-41-45.png
Original size: 734x499
Attachment 19:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown